OpenStudy (anonymous):

Martina made a block tower that was ft tall. She had a toy giraffe that was ft tall. She had another tower that was ft tall. How tall would the towers be if placed one on top of the other?

OpenStudy (anonymous):

6 11/12

OpenStudy (anonymous):

@Aveline

OpenStudy (anonymous):

@noseboy908

OpenStudy (anonymous):

@likeabossssssss

likeabossssssss (likeabossssssss):

um are there answers?

OpenStudy (anonymous):

2 1/3 ft 1 1/3 ft 3 1/4 ft

OpenStudy (anonymous):

A 4 2/3 ft B 5 7/12 ft C 6 2/3 ft D 6 11/12 ft

OpenStudy (anonymous):

D

likeabossssssss (likeabossssssss):

is there a pic of it?

OpenStudy (anonymous):

no but the fractions are in order

likeabossssssss (likeabossssssss):

um I guess u r right

OpenStudy (anonymous):

Martina made a block tower that was 2 1/3 ft tall. She had a toy giraffe that was 1 1/3 ft tall. She had another tower that was 3 1/4 ft tall. How tall would the towers be if placed one on top of the other? A. 4 2/3 ft B. 5 7/12ft C. 6 2/3 ft D. 6 11/12 ft

OpenStudy (anonymous):

@Mehek14

likeabossssssss (likeabossssssss):

one question per post plz :)

Mehek (mehek14):

@likeabossssssss same question lol

OpenStudy (anonymous):

its the same but all together

likeabossssssss (likeabossssssss):

omg my bad sorry

OpenStudy (anonymous):

its cool

Mehek (mehek14):

\(2\dfrac{1}{3}=\dfrac{28}{12}\\ 1\dfrac{1}{3}=\dfrac{16}{12}\\3\dfrac{1}{4}=\dfrac{39}{12}\\ \dfrac{28}{12}+\dfrac{16}{12}+\dfrac{39}{12}\)

OpenStudy (anonymous):

D 6 11/12

Mehek (mehek14):

correct

OpenStudy (anonymous):

thanks

Mehek (mehek14):

yw :)